OpenStudy (anonymous):

The optic nerve needs a minimum of 2.10x10^-17J of energy to trigger a series of impulses that ... How many photons of blue light (475nm) are needed? ...

OpenStudy (anonymous):

The energy of a single photon is given by E = h*c/lambda (h=planck constant, c=speed of light) E = (6.626*10^-34 J*s )*(3*10^8 m/s)/(475*10^-9 m) = 4.18484*10^-19 J/photon Divide the energy needed by the energy per photon to get the answer (2.1*10^-17 J)/(4.18484*10^-19 J/photon) = 50.18 photons (round up to 51)